THE MILLIMETERWAVE SPECTRUM OF n-BUTYL CYANIDE

\underline{MATTHIAS H. ORDU}, HOLGER S. P. MÜLLER, FRANK LEWEN, STEPHAN SCHLEMMER, I. Physikalisches Institut, Universität zu Köln, Zülpicher Str. 77, 50937 Köln, Germany; MARC NUÑEZ, and ADAM WALTERS, IRAP: Université de Toulouse, UPS-OMP, CNRS; 9 Av. colonel Roche, BP 44346, 31028 Toulouse cedex 4, France.

The rotational spectrum of n-butyl cyanide (C_{4}H_{9}CN) was measured between 75 and 130 GHz using a novel all-solid-state spectrometer with a total absorption path of 44 m. In the course of the analysis of the spectrum, about 3000 transitions were assigned and a full set of quartic centrifugal distortion parameters with some sextic and octic terms could be determined for each of the three known conformers (anti-anti, anti-gauche(methyl end) and gauche(CN end)-anti). The work was motivated by the fact that n-butyl cyanide is likely to be found in interstellar hot core environments. This is indicated by the discovery of n-propyl cyanide (C_{3}H_{7}CN), the next smaller alkyl cyanide, in the ISM^{a}. The increased accuracy of the model, which will be additionally extended by future laboratory measurements around 200 GHz, may now be employed for a prediction of the spectrum up to 300 GHz with a feasible uncertainty for astronomic line surveys. Furthermore, there are two less abundant conformers, cis-gauche-gauche and trans-gauche-gauche, which have not yet been detected in the rotational spectrum^{b}. Due to the increased sensitivity of the new spectrometer, it seems possible now for the first time to identify their sectroscopic fingerprints in the recorded data.

aA. Belloche, R. T. Garrod, H. S. P.Müller, K. M. Menten, C. Comito, and P. Schilke, Astronomy & Astrophysics 499, 215 (2009)

bR. K. Bohn, J. L. Pardus, J. August, T. Brupbacher, W. Jäger, J. Mol. Struct. 413–414, 293 (1997)